OpenStudy (anonymous):

If the point (x, √3/3) is on the unit circle, what is x? A) 2\3 B) √6\3 C) 4/9 D) √2\3

OpenStudy (amistre64):

use the pythag thrm to develop: x = sqrt(1-y^2)

OpenStudy (amistre64):

in other words: x^2 + y^2 = 1, and we know y

OpenStudy (anonymous):

So is it D?

OpenStudy (amistre64):

show me the math you did to get to D

OpenStudy (amistre64):

or, show me if: (sqrt(2)/3)^2 + (sqrt(3)/3)^2 = 1

OpenStudy (anonymous):

Okay, hold on a moment, please.

OpenStudy (anonymous):

I just plugged (sqrt(2)/3)^2 + (sqrt(3)/3)^2 = 1 into my calculator and it came out as false.

OpenStudy (amistre64):

D looks off, but it has some trick about it √(2\3) is correct, but i believe D is written as: √(2)\3 which is a different number

OpenStudy (amistre64):

\[x=\sqrt{1-\left(\frac{\sqrt{3}}{3}\right)^2}\] \[x=\sqrt{1-\frac39}\] \[x=\sqrt{\frac99-\frac39}\] \[x=\sqrt{\frac69}\] \[x=\sqrt{\frac23}~or~x=\frac{\sqrt 6}{3}\]

OpenStudy (amistre64):

so im not sure how your process took you. but if you got to sqrt(2/3) it right, but just not an option. in many cases, they dont like a sqrt on the bottom. so sqrt(6)/3 is 'more accurate'.

OpenStudy (anonymous):

Ohhhhh. I see, I think I messed up after the 3rd step! Thank you! Do you mind helping me with a few more problems? I really would appropriate it.

OpenStudy (amistre64):

i have ot be leaving soon, so i dont know if i have the time, but whatcha got?

OpenStudy (anonymous):

What is the period of the cosine and sine functions? Pi 2pi 4pi cosine and sine are not periodic I believe it's 2pi, I think I've had this problem before...but I'm not quite sure.

OpenStudy (amistre64):

cos(0) = cos(2pi) sin(0) = sin(2pi) so yes, they are periodic (they repeat themselves) at 2pi

OpenStudy (anonymous):

Thanks, okay. I have a few more...hold on a sec, please.

OpenStudy (anonymous):

Use a calculator to evaluate cot π\5. 0.561 1.376 91.185 32.142

OpenStudy (amistre64):

lolmy guess was right

OpenStudy (amistre64):

i know pi/4 has a tangent of 1, pi/5 is less than 1 by a smidgen; so id say cot has to be bigger than 1 by a smidgen ... http://www.wolframalpha.com/input/?i=cot%28pi%2F5%29

OpenStudy (anonymous):

I also got 1.376.

OpenStudy (amistre64):

my times up .... good luck with these. you seem to be doing fine so far.

OpenStudy (anonymous):

Thank you!